HUBBLE SPACE TELESCOPE DAILY REPORT #2017 PERIOD COVERED: 0000Z (UTC) 12/04/97 - 0000Z (UTC) 12/05/97 Daily Status Report as of 339/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/2/3 7811 (NICMOS Modified One Orbit Parallel) The Near Infrared Camera (Cameras Two and Three) was used to obtain "JH" imaging, and G096 and G141 spectroscopy of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.2 Completed Three Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 6431 (Determination of the Extragalactric Distance Scale) The WF/PC-2 was used to make observations of the galaxy NGC-3319 to be used in measuring the periods of Cepheid variable stars within this galaxy. This will allow for a more accurate estimate of the distances to these galaxies through the period-luminosity relation for the Cepheid stars. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Six Sets of NIC/2/3 7812 (NICMOS Modified Short One Orbit Parallel) The Near Infrared Camera (Cameras Two and Three) was used to obtain "JH" imaging of sky regions when another instrument is prime. Two dither positions are obtained for each filter. The observations were completed as scheduled, and no anomalies were reported. 1.5 Completed STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 1.6 Completed NIC/1/2 7692 (Polarizers Calibration) The Near Infrared Camera (Cameras One and Two) was used to make observations of the standard star HD283812 as part of a polarization calibration. The polarization and position angle of polarization are measured for two standard polarized point sources and one extended polarized source of known J and H band polarization. The observations were completed as scheduled, and no problems were noted. 1.7 Completed Six Sets of WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.8 Completed FOC/96 6732 (FOC Observation of the Evolution of the R-Aqr Jet) The Faint Object Camera (f/96) was used to make observations of the jet of the symbiotic binary variable star R-Agr. This data will provide a six-year baseline for studying the evolution of the binary system at a spatial resolution unattainable anywhere else for these key spectral regions. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 6905 (WF/PC-2 Cycle 6 Calibration: Internal Monitor) The WF/PC-2 was used to make several internal calibration observations including biases and internal flats as part of a Cycle 6 calibrations procedure. A variety of internal exposures are run weekly to verify and monitor the health of the camera. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD 7711 (STIS CCD G750L Fringing Flats.) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor the STIS CCD flat-field fringing in the G750L grating at it's nominal wavelength setting {7751A}. WAVECALs were also taken at the same gain and wavelength setting. Typical S/N = 100 for each integration, and gain = 4 on each. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Nine Sets of STIS/CCD 7782 (STIS Non-Scripted Parallel Archive Proposal - Half Orbit) The Space Telescope Imaging Spectrograph (CCD) was used for a non-scripted parallel archive proposal (one-half orbit). The EMC re-try described in 2.1 occurred during the seventh iteration of this proposal with no reported effect. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.13 Completed Five Sets of STIS/CCD 7601 (CCD Dark and Bias Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to conduct bias and dark monitoring. Also, observations were made to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were completed as planned, and no anomalies were reported. 1.14 Completed NIC/1/2/3 7268 (Nuclear Brightness Profiles of Merger Remnants; Do Mergers Form Ellipticals?) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the elliptical galaxy NGC-455. Many elliptical galaxies might have formed from mergers of disk galaxies. Numerical simulations yield elliptical-like remnants and observed merger remnants have R^1/4 surface brightness profiles. However, while the global properties of merger remnants and ellipticals are similar, it is unknown whether this is true for their nuclear properties as well. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed NIC/2 7331 (Near-IR Properties of the Bulges of Spiral Galaxies) The Near Infrared Camera (Camera Two) was used to make infrared observations of the bulge in spiral galaxy NGC-7742. Near-IR data at HST resolution are essential for correcting for the effect of dust and will enable us to derive robust nuclear properties for the sample. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed NIC/1/2/3 7266 (High-Resolution IR and Optical Imaging of the Fields Around a Large Sample of Z > 4.5 Quasars) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the quasar BR1033-032 in order to make high-resolution infrared and optical imaging of the fields around the quasars. This object appears to be very primitive and may represent early stages in the galaxy formation process. The observations were completed as planned, and no anomalies were reported. 1.17 Completed STIS/CCD 7783 (STIS Parallel Science) The Space Telescope Imaging Spectrograph (CCD) was used to make observations taken in parallel with the prime science. The observations were completed, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 7 At 339/060627, FGS-3 suffered a walkdown failure (scan step limit exceeded) during a single guide star acquisition for WF/PC-2 proposal 7387 (which will re described in the next report). HSTAR 6263 was written. Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: The SSR Row Isolation Test was completed at 338/1609Z. The EDAC error was isolated to array 0. Further isolation (to row) was not possible due to limitations in communications time. Additional testing will be performed at a later time. At 338/220534Z, STIS status buffer message 732 was received indicating that the EMC mechanisms (echelle blocker, mode isolation shutter, and calibration insert mechanism) had experienced a re-try. At the time the proposal in 1.12 was executing. HSTAR 6262 was written. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH